A newspaper carrier has $3.75 in change. He has three more quarters than dimes but twice as many nickels as quarters. How many coins of each type does he have?
Answer by ewatrrr(10682)


Hi
three more quarters than dimes and twice as many nickels as quarters
Let x,(x+3), 2(x+3)represent the number of dimes, quarters and nickels respectively
Question states***Note: Cents makes sense
10x+ 25(x+3)+ 5*2(x+3) = 375 CENTS
Solving for x
45x + 75 + 30 = 375
45x = 270
x = 6, the number of dimes. 9 Quarters and 18 Nickels
CHECKING our Answer***
10*6 + 25*9 + 5*18 = 60 + 225 + 90 = 375 or $3.75